OpenStudy (anonymous):

solve by completing the square: x^2-2x=15

OpenStudy (anonymous):

please show work

sam (.sam.):

x^2-2x=15 x^2-2x-15=0 (x-1)^2 (-1)-15=0 (x-1)^2-16=0 (x-1)^2=16 x-1=+-4 ---------------------------------- x-1=4 x=5 ----------------------------------- x-1=-4 x=-3

OpenStudy (anonymous):

(1) (1/2b)^2 =c1 given that b= -2, for the value of c1=[1/2(2)]2=1 following the formula a^2+bx+c1=c-c1 x^2-2x+1=15+1; given that c1 is positive. . x^2-2x+1=16 x^2-2x-15=0, i use the completing the square method, but i think there's no need for doing it, for there is already a given constant on the right side of the equation. . that's why my solution is the same with sam...